OpenStudy (anonymous):

3(x+3)-(2x+7)=9 step by step plz

OpenStudy (anonymous):

\[3(x+3)-(2x+7)=9\]\[3x+9 - 2x - 7 = 9\]\[3x-2x=9+7-9\]\[x=7\]

OpenStudy (mimi_x3):

3x + 9 -2x - 7 = 9 3x - 2x + 2 = 9 x + 2 = 9 x = 7